# GEMeX-ThinkVG: Towards Thinking with Visual Grounding in Medical VQA via Reinforcement Learning

### v1

@@ Abstract
Medical visual question answering aims to support clinical decision-making by enabling models to answer natural language questions based on medical images. While recent advances in multi-modal learning have significantly improved performance, current methods still suffer from limited answer reliability and poor interpretability, impairing the ability of clinicians and patients to understand and trust model-generated answers. To address this, this work first proposes a Thinking with Visual Grounding (ThinkVG) dataset wherein the answer generation is decomposed into intermediate reasoning steps that explicitly ground relevant visual regions of the medical image, thereby providing fine-grained explainability. Furthermore, we introduce a novel verifiable reward mechanism for reinforcement learning to guide post-training, improving the alignment between the model’s reasoning process and its final answer. Remarkably, our method achieves comparable performance using only one-eighth of the training data, demonstrating the efficiency and effectiveness of the proposal. The dataset is available at https://huggingface.co/datasets/BoKelvin/GEMeX-ThinkVG.
1 Introduction
Figure 1: Interpretability comparison: (Top) Traditional answer with grounding provided by GEMeX: separate textual and visual prompts; (Bottom) Our integrated, more detailed thinking process for answer generation, which explicitly grounds evidence in specific regions (shown in the colored texts) of the medical image, i.e., anatomical areas that support inferences, thereby enhancing the understanding of questions and answers.
@@ 1 Introduction
Medical Visual Question Answering (VQA) has emerged as a promising paradigm for supporting clinical decision-making by enabling models to answer natural language questions based on medical images [16]. Recent advances in multi-modal learning have led to significant performance improvements across a range of Medical VQA benchmarks [19, 13, 8, 6], especially in the era of large vision language models [17, 34, 3, 36]. However, despite these gains, existing methods remain limited in their ability to provide interpretable and trustworthy responses—an essential requirement for real-world deployment in clinical settings. In particular, most current approaches generate answers without revealing the underlying reasoning process or the specific image regions that inform their decisions. This lack of transparency undermines user trust and hinders the integration of such systems into clinical workflows.
@@
A recent work GEMeX [21] introduced a new dataset in which each VQA triplet is paired with a textual reason and a corresponding visual region as an explanation, bridging this gap to some extent. However, the textual and visual prompts are independent and lack coherence. Moreover, the textual reason merely serves as an explanation of the answer rather than providing guidance on how to solve the question based on the image. Therefore, based on GEMeX, we propose a new dataset GEMeX-ThinkVG that enhances interpretability in Medical VQA through introducing a Thinking with Visual Grounding mechanism, as shown in Figure 1. Different from the traditional chain-of-thought reasoning [35] and recent approaches [24, 15] that rely solely on reinforcement learning to incentivize the thinking process for solving multimodal tasks, our ThinkVG explicitly decomposes the answer generation process into a sequence of intermediate steps, along with grounding corresponding visual evidence within the medical image. By linking reasoning steps to spatially localized regions, we offer clinicians a fine-grained explanation of how the model arrives at its conclusions.
@@
While leveraging the GEMeX-ThinkVG dataset for Medical VQA enables the modeling of explicit thinking, our empirical findings indicate that supervised fine-tuning (SFT) alone is insufficient to fully unleash the reasoning capabilities of the large vision language models (LVLMs). To address this, we adopt a reinforcement learning (RL) framework for post-training that builds upon the SFT-tuned model. It introduces a novel verifiable reward mechanism to incentivize alignment between the reasoning trajectory and relevant visual and textual cues, thereby enhancing both the accuracy of intermediate thinking steps and final answer generation. In contrast to existing methods that require large-scale datasets, our approach is data-efficient, achieving comparable performance using only one-eighth of the full training data.
@@
  • We introduce GEMeX-ThinkVG, a Thinking with Visual Grounding dataset for Medical VQA, which provides step-by-step reasoning grounded in specific anatomical regions of medical images, offering better interpretability for understanding the questions and answers.
  • We perform supervised fine-tuning on the LVLM using our GEMeX-ThinkVG dataset, and further apply reinforcement learning with a novel verifiable reward mechanism to incentivize reasoning abilities, enabling the generation of more accurate thinking paths and final answers.
@@
2 Related Work
2.1 Medical VQA Dataset
@@
Recent years have witnessed significant progress in medical visual question answering (VQA) through the development of specialized datasets targeting diverse clinical challenges. VQA-RAD [16] established the foundation with 3,000 question-answer pairs focused on radiology images. SLAKE [20] expanded the scope with over 14,000 manually annotated QA pairs spanning CT, MRI, and X-ray modalities. OmniMedVQA [8] further broadened coverage across multiple body regions and imaging modalities to enhance model generalization capabilities. For specialized tasks, PathVQA [6] offers 32,000 QAs on histopathology images. MIMIC-Diff-VQA [7] emphasizes differential diagnosis between paired X-rays. GEMeX [21] enhances VQA with both textual and visual explanations, facilitating better understanding of the answers for patients. However, none of them incorporate a reasoning process for problem-solving, which hinders patients and junior doctors from fully understanding the questions and answers. To fill this gap, we propose a thinking with visual grounding dataset, GEMeX-ThinkVG, in this work.
2.2 Reinforcement Learning
@@
Recently, the growing popularity of reasoning-focused large language models such as GPT-o1 [10] and DeepSeek-R1 [5] has brought increasing attention to the underlying mechanisms of reinforcement learning (RL). This interest has spurred extensive research and led to remarkable advancements in tasks such as mathematics [23, 32], code generation [12, 33], and hallucination detection [37, 30]. A notable contribution is the Group Relative Policy Optimization (GRPO) algorithm proposed in DeepSeekMath [28], which significantly simplifies the training process. Unlike traditional RL algorithms like PPO [12] that require a critic model to evaluate policy performance, GRPO compares groups of candidate responses directly, eliminating the need for an additional critic. GRPO’s simplicity has inspired further exploration into the reasoning capabilities of large vision-language models (LVLMs) for multimodal and visual tasks [9, 22, 29]. In the medical domain, GRPO has also been applied to incentivize the reasoning abilities of LVLMs for tasks such as out-of-distribution detection [24] and visual question answering [15]. Unlike these works that merely elicit textual descriptions of visual content, we propose a visually grounded reasoning paradigm that provides improved interpretability.
3 Construction of GEMeX-ThinkVG
@@
Figure 2: An illustration of the proposed uncertainty-driven multi-agent framework for generating the GEMeX-ThinkVG dataset. The left (orange) section depicts the data generation process, while the right (blue) section shows the uncertainty-driven quality assurance pipeline, i.e., the “not sure” option prompts the agent to engage in deeper reflection and more deliberate decision-making.
3.1 Task Definition
@@
Unlike conventional chain-of-thought approaches used in general tasks such as VQA [31], which typically generate only textual rationales before outputting the final answer, we propose a multimodal-level thinking process with visual grounding tailored for medical VQA. As illustrated in Figure 1, it decomposes the reasoning process into several steps with visual attention to specific image regions, aiming to deliver more detailed explanations and improved interpretability for patients and junior doctors. Formally, given a question $q$ for image $v$, we hope to provide a multi-modal thinking process $T$ involving visual grounding to attain answer $a$ via
$T=\left(\{r_{i}\}_{i=1}^{N_{r}},\{t_{j}\}_{j=1}^{N_{t}}\right),$ (1)
@@
3.2 An Uncertainty-Driven Multi-Agent Framework
As illustrated in Figure 2, we construct the ThinkVG dataset based on GEMeX [21], the largest Medical VQA dataset that provides detailed visual and textual explanations for each VQA triplet. To ensure the high quality of ThinkVG, we introduce an uncertainty-driven multi-agent framework, centered around DeepSeek-R1 [5], which also coordinates multiple agents and leverages an uncertainty mechanism to enhance the accuracy of the generated data.
3.2.1 Data Generation
In Stage 1, we primarily use DeepSeek-R1 to generate the thinking process for each VQA triplet due to its outstanding reasoning capabilities. However, as DeepSeek-R1 is a text-based LLM and lacks the ability to directly perceive visual inputs, we utilize the grounded reports from GEMeX as an intermediary. In GEMeX’s reports, each sentence is aligned with a specific anatomical region, enabling DeepSeek-R1 to simulate a multimodal reasoning process required for solving the tasks. To optimize the prompt design, we randomly sample 50 questions from each question type (e.g., open-ended) in the training set—resulting in 200 questions in total—to evaluate performance and guide prompt iteration. The final optimized prompt is presented in Table 1.
@@ 3.2.1 Data Generation
Table 1: Given a question-answer pair, our tailored prompt guides DeepSeek-R1 to generate ThinkVG data.
3.2.2 Uncertainty-driven Quality Assurance
Although manual verification and prompt modification are applied, the generated data still suffers from quality issues. To address this, we design a pipeline which contains three steps (Stage 2 to Stage 4) to ensure quality: (1) We first introduce an uncertainty-aware mechanism that enables DeepSeek-R1 to self-reflect and evaluate the reliability of its own outputs. Specifically, after generating the ThinkVG response, we prompt DeepSeek-R1 with the question: “Do you trust your thinking process? A. Yes B. No C. Not sure”. This self-reflection with uncertainty choice “Not sure” allows the model to carefully reassess its previous reasoning, which has been shown to improve output accuracy [2]. We retain only those ThinkVG responses for which the model answers “A. Yes.”; (2) To avoid bias caused by relying on a single LLM for both data generation and evaluation [25], we introduce additional agents to assess the quality of the generated data. As shown in Stage 3, we employ three additional agents—Qwen-2.5-72B [26], OpenBioLLM-70B [^1] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B, and GPT-4o [1]—as external reviewers. These agents are given grounded reports, questions, and answers to assess the accuracy of the thinking process. Similar to the self-reflection stage, we introduce an “uncertain” option to provide the agents with greater flexibility for consideration and decision-making. This process helps eliminate incorrect reasoning paths generated by DeepSeek-R1 based on its internal knowledge. Then, we adopt a voting mechanism: the ThinkVG is retained only if all three agents unanimously select “A. Correct” option; (3) Finally, we apply post-processing to eliminate ThinkVG outputs with significant errors, such as incorrect location information or instances that quote the report or findings directly, rather than generating content from an observational perspective.
3.3 Radiologist Validation
After auto-generation, a radiologist randomly samples 200 examples from each of the four question types in the dataset for manual inspection. Cross-validation reveals no clinically inconsistent reasoning chains, demonstrating the effectiveness of the proposed framework. In total, we generate 206,071 ThinkVG instances for 21,994 images, all drawn from the GEMeX training set. The detailed distribution is presented in Table 2.
Figure 3: An example from the SFT and RFT stages: after SFT, the model learns to think carefully and incorporates grounding to specific pathological regions for visual evidence. During the RFT stage, since both the answer and the grounded regions are correct, the two reward scores are each 1.0.
@@ 3.3 Radiologist Validation
62,712 37,340 53,905 52,114
Table 2: Distribution of generated ThinkVG data across the four question types in GEMeX.
4 Fine-tuning on GEMeX-ThinkVG: Learning to Think with Visual Grounding in Medical VQA
4.1 Supervised Fine-tuning
After generating the ThinkVG data, we first train the large vision language model (LVLM) using Supervised Fine-tuning (SFT) to equip it with task-aware thinking capabilities. To generate the outputs required by GEMeX—which include both the final answer and its corresponding visual location (omitting the reasons, as we provide a more detailed preceding thinking path)—the model must produce structured outputs in two distinct parts. First, it should output a thinking trace that incorporates explicit visual grounding, enclosed within the <think></think> tags. Second, it should deliver a concise final response that includes the predicted “answer” and the summarized “visual location”, wrapped within a designated XML-like structure: <response> <answer></answer> <location></location></response>. This format ensures both interpretability and alignment with the GEMeX evaluation requirements. An illustrative example is shown in Figure 3.
4.2 Reinforcement Fine-tuning
Although SFT can significantly improve thinking capabilities, we have observed in practice that: (1) due to the complexity of the task, the model sometimes generates incorrect thinking process after SFT, for example, incorrect analysis of images, which may be because SFT is better at helping LVLMs memorize rather than truly understand  [4]; (2) inaccurate region localization within the thinking path can propagate errors to the final answer, ultimately leading to degraded performance. Therefore, we adopt Reinforcement Fine-tuning (RFT), specifically through Group Relative Policy Optimization (GRPO) [28], as a post-training strategy to further activate and enhance the model’s thinking capabilities with a newly proposed reward mechanism.
4.2.1 Group Relative Policy Optimization

### v2

# GEMeX-RMCoT: An Enhanced Med-VQA Dataset for Region-Aware Multimodal Chain-of-Thought Reasoning

@@ Abstract.
Medical visual question answering aims to support clinical decision-making by enabling models to answer natural language questions based on medical images. While recent advances in multi-modal learning have significantly improved performance, current methods still suffer from limited answer reliability and poor interpretability, impairing the ability of clinicians and patients to understand and trust model outputs. To address these limitations, this work first proposes a Region-Aware Multimodal Chain-of-Thought (RMCoT) dataset, in which the process of producing an answer is preceded by a sequence of intermediate reasoning steps that explicitly ground relevant visual regions of the medical image, thereby providing fine-grained explainability. Furthermore, we introduce a novel verifiable reward mechanism for reinforcement learning to guide post-training, improving the alignment between the model’s reasoning process and its final answer. Remarkably, our method achieves comparable performance using only one-eighth of the training data, demonstrating the efficiency and effectiveness of the proposal. The dataset is available at https://www.med-vqa.com/GEMeX/.
Large vision language models, Medical visual question answering, Explainable and trustworthy AI, Chest X-ray dataset
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3758277ccs: Computing methodologies Natural language generationccs: Applied computing Health care information systemsccs: Computing methodologies Computer vision tasks
1. Introduction
Figure 1. (Top) Traditional answer with explainability provided by GEMeX: separate textual and visual prompts;
[…]
@@ 1. Introduction
Medical Visual Question Answering (Med-VQA) has emerged as a promising paradigm for supporting clinical decision-making by enabling models to answer natural language questions based on medical images (Lau et al., 2018). Recent advances in multi-modal learning have led to significant performance improvements across a range of Med-VQA benchmarks (Liu et al., 2021a; Khare et al., 2021; Hu et al., 2024; He et al., 2020), especially in the era of large vision language models (Li et al., 2024; Zhang et al., 2023a; Chen et al., 2024; Sellergren et al., 2025; Cui et al., 2025; Zou et al., 2025). However, despite these gains, existing methods remain limited in their ability to provide interpretable and trustworthy responses—an essential requirement for real-world deployment in clinical settings. In particular, most current approaches generate answers without revealing the underlying reasoning process or the specific image regions that inform their decisions. This lack of transparency undermines user trust and hinders the integration of such systems into clinical workflows.
@@
Recently, GEMeX (Liu et al., 2024) introduced a new dataset where each VQA triplet is paired with a textual reason and a corresponding visual region as an explanation, bridging this gap to some extent. However, the textual and visual prompts are independent and lack coherence. Moreover, the textual reason merely serves as an explanation of the answer rather than providing guidance on how to solve the question based on the image. Therefore, based on it, we propose a new dataset GEMeX-RMCoT that enhances interpretability through introducing a Region-Aware Multimodal Chain-of-Thought reasoning mechanism, as shown in Figure 1. Different from the traditional CoT (Zhang et al., 2023b) and recent approaches (Pan et al., 2025; Lai et al., 2025) that rely solely on reinforcement learning to incentivize the thinking process, our RMCoT explicitly decomposes answer generation into intermediate steps, along with grounding corresponding visual evidence within the medical image. By linking reasoning steps to spatially localized regions (Su et al., 2025b), we offer clinicians and patients a fine-grained explanation of how the model arrives at its conclusions.
@@
While the GEMeX-RMCoT dataset enables the modeling of explicit thinking, our empirical findings indicate supervised fine-tuning (SFT) alone is insufficient to fully unleash the reasoning capabilities of the large vision language models (LVLMs). To address this, we employ reinforcement learning for post-training that builds upon the SFT-tuned model. It introduces a novel verifiable reward mechanism to incentivize alignment between the reasoning trajectory and relevant visual and textual cues, thereby enhancing both the accuracy of intermediate thinking steps and final answer generation. In contrast to existing methods requiring large-scale datasets, our approach is data-efficient, achieving comparable performance using only one-eighth of the full training data.
@@
  • We introduce GEMeX-RMCoT, a region-aware multimodal chain-of-thought dataset for Med-VQA, which provides step-by-step reasoning explicitly grounded in specific anatomical regions of medical images, offering better interpretability for understanding the questions and answers.
  • We perform supervised fine-tuning on the LVLM using our GEMeX-RMCoT dataset, and further apply reinforcement learning with a novel verifiable reward mechanism to incentivize reasoning abilities, enabling the generation of more accurate thinking paths and final answers.
@@
2. Related Work
2.1. Medical VQA Dataset
@@
Recent years have witnessed significant progress in medical visual question answering through the development of specialized datasets targeting diverse clinical challenges. VQA-RAD (Lau et al., 2018) established the foundation with 3,000 question-answer pairs focused on radiology images. SLAKE (Liu et al., 2021b) expanded the scope with over 14,000 manually annotated QA pairs spanning CT, MRI, and X-ray modalities. OmniMedVQA (Hu et al., 2024) further broadened coverage across multiple body regions and imaging modalities to enhance model generalization capabilities. For specialized tasks, PathVQA (He et al., 2020) offers 32,000 QAs on histopathology images. MIMIC-Diff-VQA (Hu et al., 2023) emphasizes differential diagnosis between paired X-rays. GEMeX (Liu et al., 2024) enhances VQA with both textual and visual explanations, facilitating a better understanding of the answers. However, none of them incorporates a reasoning process for problem-solving, which hinders patients and junior doctors from fully understanding the questions and answers. To fill this gap, we propose a region-aware multimodal CoT reasoning dataset, GEMeX-RMCoT, in this work.
2.2. Reinforcement Learning
@@
Recently, the growing popularity of reasoning-focused large language models such as GPT-o1 (Jaech et al., 2024) and DeepSeek-R1 (Guo et al., 2025) has drawn increasing attention to the underlying mechanisms of reinforcement learning (RL). This interest has led to remarkable advancements in tasks such as mathematics (Luong et al., 2024; Ying et al., 2024), code generation (Jiao et al., 2024; Zhang et al., 2024), hallucination detection (Zhao et al., 2023; Sun et al., 2023), and interdisciplinary research (Zhao et al., 2025; Zhou et al., 2025). A notable contribution is the Group Relative Policy Optimization (GRPO) algorithm proposed in DeepSeekMath (Shao et al., 2024), which significantly simplifies the training process. Unlike traditional RL algorithms like PPO (Jiao et al., 2024) that require a critic model to evaluate policy performance, GRPO compares groups of candidate responses directly, eliminating the need for an additional critic. GRPO’s simplicity has inspired further exploration into the reasoning capabilities of LVLMs for multimodal and visual tasks (Huang et al., 2025; Liu et al., 2025; Su et al., 2025a). In the medical domain, GRPO has also been applied to incentivize the reasoning abilities of LVLMs for tasks such as out-of-distribution detection (Pan et al., 2025) and visual question answering (Lai et al., 2025). Unlike these works that merely elicit textual descriptions of visual content, we propose a visually grounded reasoning paradigm that enhances interpretability and strengthens evidential support.
3. Construction of GEMeX-RMCoT
@@
Figure 2. Illustration of the proposed uncertainty-driven multi-agent framework for generating the RMCoT dataset. The left (orange) section depicts the data generation process, while the right (blue) section shows the uncertainty-driven quality assurance pipeline, i.e., the “not sure” option prompts a deeper reflection and more deliberate decision-making.
3.1. Task Definition
@@
Unlike conventional chain-of-thought approaches used in general tasks such as VQA (Wang et al., 2025), which typically generate only textual rationales before outputting the final answer, we propose a multimodal-level thinking process with visual grounding tailored for Med-VQA. As illustrated in Figure 1, it decomposes the reasoning process into several steps with visual attention to specific image regions, aiming to deliver more detailed explanations and improved interpretability for patients and junior doctors. Formally, given a question $q$ for image $v$, we hope to provide a multi-modal thinking process $T$ involving visual grounding to attain answer $a$ via
(1) $T=\left(\{r_{i}\}_{i=1}^{N_{r}},\{t_{j}\}_{j=1}^{N_{t}}\right),$
@@
3.2. An Uncertainty-Driven Multi-Agent Framework
As illustrated in Figure 2, we construct the RMCoT dataset based on GEMeX (Liu et al., 2024), the largest Med-VQA dataset that provides detailed visual and textual explanations for each VQA triplet. To ensure the high quality of RMCoT, we introduce an uncertainty-driven multi-agent framework, centered around DeepSeek-R1 (Guo et al., 2025), which also coordinates multiple agents and leverages an uncertainty mechanism to enhance the accuracy of the generated data.
3.2.1. Data Generation
In Stage ①, we primarily use DeepSeek-R1 to generate the thinking process for each VQA triplet due to its outstanding reasoning capabilities. However, as DeepSeek-R1 is a text-based LLM and lacks the ability to directly perceive visual inputs, we utilize the grounded reports from GEMeX as an intermediary. In GEMeX’s reports, each sentence is aligned with a specific anatomical region, enabling DeepSeek-R1 to simulate a multimodal reasoning process required for solving the tasks. To optimize the prompt design, we randomly sample 50 questions from each question type (e.g., open-ended) in the training set—resulting in 200 questions in total—to evaluate performance and guide prompt iteration. The final optimized prompt is presented in Table 1.
@@ 3.2.1. Data Generation
Table 1. Given a question-answer pair, our tailored prompt guides DeepSeek-R1 to generate RMCoT data.
3.2.2. Uncertainty-driven Quality Assurance
Although manual verification and prompt modification are applied, the generated data still suffers from quality issues. To address this, we design a pipeline which contains three steps (Stage ② to Stage ④) to ensure quality: (1) We first introduce an uncertainty-aware mechanism that enables DeepSeek-R1 to self-reflect and evaluate the reliability of its own outputs. Specifically, after generating the RMCoT response, we prompt DeepSeek-R1 with the question: “Do you trust your thinking process? A. Yes B. No C. Not sure”. This self-reflection with uncertainty choice “Not sure” allows the model to carefully reassess its previous reasoning, which has been shown to improve output accuracy (Chen and Mueller, 2023). We retain only those RMCoT responses for which the model answers “A. Yes.”; (2) To avoid bias caused by relying on a single LLM for both data generation and evaluation (Panickssery et al., 2024), we introduce additional agents to assess the quality of the generated data. As shown in Stage ③, we employ three additional agents—Qwen-2.5-72B (Qwen et al., 2024), OpenBioLLM-70B [^1] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B, and GPT-4o (Achiam et al., 2023)—as external reviewers. These agents are given grounded reports, questions, and answers to assess the accuracy of the thinking process. Similar to the self-reflection stage, we introduce an “uncertain” option to provide the agents with greater flexibility for consideration and decision-making. This process helps eliminate incorrect reasoning paths generated by DeepSeek-R1 based on its internal knowledge. Then, we adopt a voting mechanism: the RMCoT is retained only if all three agents unanimously select “A. Correct” option; (3) Finally, we apply post-processing to eliminate RMCoT outputs with significant errors, such as incorrect location information or instances that quote the report or findings directly, rather than generating content from an observational perspective.
3.3. Radiologist Validation
After auto-generation, a radiologist randomly samples 200 examples from each of the four question types in the dataset for manual inspection. Cross-validation reveals no clinically inconsistent reasoning chains, demonstrating the effectiveness of the proposed framework. In total, we generate 202,384 RMCoT instances for 21,994 images, all drawn from the GEMeX training set. The detailed distribution is presented in Table 2.
Figure 3. An example from the SFT and RFT stages: after SFT, the model learns to think carefully and incorporates grounding to specific pathological regions for visual evidence. During the RFT stage, since both the answer and the grounded regions are correct, the two reward scores are each 1.0.
@@ 3.3. Radiologist Validation
61,240 36,432 52,850 51,862
Table 2. Distribution of generated RMCoT data across the four question types in GEMeX.
4. Fine-tuning on GEMeX-RMCoT: Learning to Think with Visual Grounding in Med-VQA
4.1. Supervised Fine-tuning
After generating the RMCoT data, we first train the large vision language model (LVLM) using Supervised Fine-tuning (SFT) to equip it with task-aware thinking capabilities. To generate the outputs required by GEMeX—which include both the final answer and its corresponding visual location (omitting the reasons, as we provide a more detailed preceding thinking path)—the model must produce structured outputs in two distinct parts. First, it should output a thinking trace that incorporates explicit visual grounding, enclosed within the <think></think> tags. Second, it should deliver a concise final response that includes the predicted “answer” and the summarized “visual location”, wrapped within a designated XML-like structure: <response> <answer></answer> <location></location></response>. This format ensures both interpretability and alignment with the GEMeX evaluation requirements. An illustrative example is shown in Figure 3.
4.2. Reinforcement Fine-tuning
Although SFT can significantly improve thinking capabilities, we have observed in practice that: (1) due to the complexity of the task, the model sometimes generates incorrect thinking process after SFT, for example, incorrect analysis of images, which may be because SFT is better at helping LVLMs memorize rather than truly understand  (Chu et al., 2025); (2) inaccurate region localization within the thinking path can propagate errors to the final answer, ultimately leading to degraded performance. Therefore, we adopt Reinforcement Fine-tuning (RFT), specifically through Group Relative Policy Optimization (GRPO) (Shao et al., 2024), as a post-training strategy to further activate and enhance the model’s thinking capabilities with a newly proposed reward mechanism.
4.2.1. Group Relative Policy Optimization